OpenStudy (anonymous):

General admission tickets to the fair cost $3.50 per person. Ride passes cost an additional $5.50 per person. Parking cost $6 for the family. The total cost for ride passes and parking was $51. How many people in the family attended the fair?

OpenStudy (vinicius1):

Not sure if I got it right: since it costs 3.5 per person and an additional fee of 5.5 per person, I'm just going to assume its 9 per person (5.5+3.5). Parking cost is 6 for the family and the total was 51.

OpenStudy (vinicius1):

So, x = number of people in the family 9x+6=51 9x=51-6 x=45/9 x=5